OpenStudy (anonymous):

Find the volume of a cone whose diameter is 15 mm and height is 17 mm. (round answer to two decimal places)

OpenStudy (anonymous):

since the volume of a cone is pir^2h/3. u are given the diameter her. so what would be the radius?

OpenStudy (jdoe0001):

 \(\bf \textit{volume of a cone}=\cfrac{\pi {\color{brown}{ r}}^2 {\color{blue}{ h}}}{3}\qquad {\color{blue}{ h}}=height\qquad {\color{brown}{ r}}=radius=\frac{diameter}{2}\)

OpenStudy (anonymous):

so the diameter would be 15/2= 7.5

OpenStudy (anonymous):

I meant the radius

OpenStudy (anonymous):

so 7.5^2=56.25

OpenStudy (anonymous):

56.25pi(17)/3=?

OpenStudy (anonymous):

either 318.75 pi or 1001.38 depending if they want ur answer with pi or multiplied out with pi

OpenStudy (anonymous):

@Marina17 are any of those the answers? if its multiple choice

OpenStudy (anonymous):

it isn't multiple choice, I have to write it